find the value of k which the equation: 3xsquared-kx+7=0 will have two real solutions
please help if you can thanks

Answer by checkley77(12844)

3x^2-kx+7=0
3x^2-22x+7=0
(3x-1)(x-7)=0
3x-1=0
3x=1
x=1/3 ans.
x-7=0
x=7 ans.
OR:
3x^2-10x+7=0
(3x-7)(x-1)=0
3x-7=0
3x=7
x=7/3 ans
x-1=0
x=1 ans.